OpenStudy (anonymous):

If 3^2y = 16 then 3^y = ?

OpenStudy (anonymous):

take sqrt bouth sides

OpenStudy (anonymous):

\(\large 3^{2y}=(3^y)^2=16 \) so 3^y = 4

OpenStudy (anonymous):

Thank you very much! I never thought of doing that. It's been a while since I touched exponential rules.